OpenStudy (anonymous):

it takes 80 j of work to stretch a spring 0.5 m from its equilibrium position. how much work is needed to stretch it an additional 0.5m a)240 j b)300 j c) 400j d)500j

OpenStudy (anonymous):

HI :D

ganeshie8 (ganeshie8):

\(\large W = -\frac{1}{2} k x^2 \) plugin x = 0.5 and find \(k\) first

ganeshie8 (ganeshie8):

\(\large 80 = -\frac{1}{2} k 0.5^2\) solve \(k\)

OpenStudy (anonymous):

=12.64 or =4square root of 10

ganeshie8 (ganeshie8):

nope, try again :)

ganeshie8 (ganeshie8):

\(80 = -\frac{1}{2} k 0.5^2 \) multiply 2 both sides \(160 = -k 0.5^2\) divide 0.5^2 both sides \(160/0.5^2 = -k\) \(k = -640\)

ganeshie8 (ganeshie8):

So, \(W = \frac{1}{2}640\times x^2 = 320 \times x^2\) plugin x = 1 to find the total work done in streteching from equilibrium to 1 meter \(W = 320 \times 1^2 = 320\) since we wanto knw the work done in moving from 0.5 to 1m subtract 80J from this :- \(320 - 80 = 240J\)

OpenStudy (anonymous):

thanks so much this was a lot of help!

ganeshie8 (ganeshie8):

np :)